A 10% coupon bond that pays coupons semiannually, matures in 5 years, and its yield-to-maturity is 8%. The bond has a face value of $1000. What is the intrinsic value of the bond today?
Answer #1

Intrinsic Value of the bond is present value of all future cash outflows in the form of interest and redemption amount.

Given information

Face Value 1000 , YTM (Yield to maturity) = 4 % (i.e. semi annually)

Interest is 1000*5%= 50 ( Semi annually ), Period 5 Years

Calculation of Intrinsic Value of the bond

= Present Value of interest @ 4 % + Present Value of redemption amount @ 4%

= 50 * 8.11 + 1000 * 0.68

= 405.50 + 680

= 1085.50

Note : It is assumed that redemption will be at par.